OpenStudy (icalibear):

Write an equation that relates x and y? (table below) x | y ---- 1 | 1 2 | 3 3 | 7 4 | 15 5 | 31 6 | 63 7 | 127 Write an equation that relates x and y. Use your equation to find the value of y when x = 15. Please help

OpenStudy (paxpolaris):

it is not linear progression the difference in the consecutive terms keeps doubling

OpenStudy (paxpolaris):

\[\Large y= 2^x-1\]

OpenStudy (icalibear):

ok, let me try that

OpenStudy (icalibear):

that wouldn't work for the first set because 1 x 1 = 1, - 1 = 0.. don't I need to get 3

OpenStudy (paxpolaris):

1=2^1 -1 = 2 - 1

OpenStudy (icalibear):

still doesn't equal 3 though

OpenStudy (paxpolaris):

3= 2^2 -1 7= 2^3 -1 15= 2^4 -1 ...

OpenStudy (icalibear):

ohok

OpenStudy (icalibear):

so y = x^2 +1

OpenStudy (icalibear):

still doesn't work

OpenStudy (paxpolaris):

nooo, y= 2^x -1

OpenStudy (paxpolaris):

when x=5 y=2^5 -1 = 2*2*2*2*2 -1

OpenStudy (icalibear):

no because for the third on, 7, 7 = 3^2 - 1 = 8. it doesnt work

OpenStudy (icalibear):

one*

OpenStudy (paxpolaris):

when x = 3 y = 2^3 -1 = 2*2*2 -1

OpenStudy (paxpolaris):

2^1 = 2 2^2 = 2*2 2^3 = 2*2*2 2^4 = 2*2*2*2 2^5 = 2*2*2*2*2 2^6 = 2*2*2*2*2*2

OpenStudy (icalibear):

yes but that doesn't work for all of them

OpenStudy (icalibear):

 etc..